OpenStudy (anonymous):

There are 250 students taking either band or chorus. If there are 180 students taking band and 60 students in both band and chorus, how many students are in just chorus?

OpenStudy (anonymous):

250-180= 70..unless you count the people in both band and cgourus

OpenStudy (anonymous):

To re-word this a bit to make it clearer: There are 250 students in total. There are 180 students in band, and 60 of those students are also in chorus. Therefore, there are (180 - 60 = 120) students are are in band ONLY, and the rest are in chorus. 250 - 120 = 130 students in chorus.